# The Eye as a Window to Systemic Health: A Survey of Retinal Imaging from Classical Techniques to Oculomics

### v1

@@ Abstract
The unique vascularized anatomy of the human eye, encased in the retina, provides an opportunity to act as a window for human health. The retinal structure assists in assessing the early detection, monitoring of disease progression and intervention for both ocular and non-ocular diseases. The advancement in imaging technology leveraging Artificial Intelligence has seized this opportunity to bridge the gap between the eye and human health. This track paves the way for unveiling systemic health insight from the ocular system and surrogating non-invasive markers for timely intervention and identification. The new frontiers of oculomics in ophthalmology cover both ocular and systemic diseases, and getting more attention to explore them. In this survey paper, we explore the evolution of retinal imaging techniques, the dire need for the integration of AI-driven analysis, and the shift of retinal imaging from classical techniques to oculomics. We also discuss some hurdles that may be faced in the progression of oculomics, highlighting the research gaps and future directions.
Index Terms:
Retinal Imaging, Retinal Vasculature, Oculomics, Systemic Diseases, Artificial Intelligence, Deep Learning, Retinal Biomarkers, Phenotypes, Genotypes
@@ I Introduction
The retina, the crucial light-sensitive tissue at the back of the eye, is a treasure trove of insights for ophthalmologists, clinicians, and technicians. It plays a pivotal role in the early prediction and intervention of both ocular and non-ocular disorders. Retinal imaging (RI), a key interpretation, encompasses a variety of techniques designed to visualise and analyse the retina [1]. Fundus photography (FP) captures detailed, subtle changes in the ocular system, while optical coherence tomography (OCT) provides a cross-sectional view, and OCT angiography (OCTA) visualizes blood flow. These modalities, serving as vital diagnostic tools, enable healthcare professionals to examine the intricate structure of the retina, including the macula, optic disc, and retinal vasculature. This non-invasive assessment allows for the early detection and monitoring of ocular diseases like diabetic retinopathy (DR), glaucoma (GL), macular degeneration (MD), and other related diseases that can lead to significant vision impairments if left unattended [2].
@@
Beyond ocular health, RI also provides insight into detecting systemic diseases. The retina’s unique accessibility and the visibility of its microvasculature make it a valuable indicator of overall health. For example, changes observed in the blood vessels and layers of the retina can trigger cardiovascular (CV) and cerebrovascular diseases. That is why the eye serves as a window of overall health in the association of ocular and systemic health. High-resolution imaging can further reveal fine-grained features and act as a facilitator for timely intervention and personalised care. This contributes to improved patient outcomes and a better understanding of the interrelation between bodily systems [3, 4].
@@
Oculomics studies ocular biomarkers, using advanced imaging technologies to probe the eye in relation to other organ systems through shared physiological functions, embryological origins, and anatomical structures. It aims to establish correlations between these markers and systemic diseases. Growing evidence indicates that oculomics is significant in identifying clinical manifestations in ocular markers, such as retinal vessel patterns and neural layer thickness, which can reveal the presence of systemic diseases such as diabetes, CV diseases (CVDs), neurological issues, and stress-related disorders. This field is driven by three significant advances, including the widespread availability of high-resolution ophthalmic imaging, the emergence of big data, and the development of automated artificial intelligence (AI) software. Together, it has the potential to revolutionise disease management by providing a comprehensive and accessible approach to monitoring and maintaining human health [5, 6].
@@
AI is at its peak in the era of big data and parallel processing. The AI model is based on recorded data, and disciplines with data can easily participate in it. In the current epoch, it is hard to conclude that there may be fields in academia and industry that can thrive without data. Therefore, no discipline will be empty-handed when discussing data. AI models are now prevalent and deployed in every aspect of life [7]. Both intrinsic and extrinsic deployments have positively impacted our society, especially in socioeconomic and socio-health aspects. Renowned computer scientist Dr. Knuth revealed that biology has at least 500 years of exciting problems to explore [8]. This perspective encourages the computing and biomedical research communities to address challenges in life science collaboratively.
@@
AI is transforming RI and ocular studies by automating and enhancing the analysis of complex visual data. AI algorithms, especially deep learning (DL), excel in identifying subtle patterns and abnormalities within retinal images that human observers might overlook. This automation significantly speeds up the screening process for ocular and non-ocular diseases, allowing for earlier detection and intervention. In addition, AI can enable the development of predictive models by evaluating the vast data set, encompassing the assessment of the individual patient’s risk factors and anticipating disease progression, paving the way for personalised medicine in ophthalmology [9, 10, 11].
@@
Eye diagnostic research is shifting from its classical origin to the current AI software tools. Now, it’s beyond the classical methods, such as preprocessing, model training, and limited clinically relevant output. These limited circumstances hinder the quantification of retinal structures at the minute and micro levels to address the objective measurements, which are crucial for the efficacy of effective treatments and the monitoring of disease progression in the growing field of oculomics. Likewise, for meaningful research, high-quality and large health datasets comprising rich annotated data are crucial, which advances the role of statistical, computational, and AI-driven techniques [12, 13]. Furthermore, this framework can be further nourished by integrating retinal imaging with clinical information, namely electronic health records (EHRS) and genetic data, including diverse demographic and ethnic data. However, translating this vision into practice remains a considerable challenge. Existing datasets often lack the necessary integration across epidemiological, clinical, and imaging domains, limiting the development of a unified infrastructure supporting oculomics research and clinical application [4, 14, 15].
@@
Despite the increasing availability of large-scale datasets, significant limitations impede their practical use in oculomics. Even though there are several Retinal imaging repositories, including those of Moorfields Eye Hospital [16], UK Biobank [17], and the 10k Cohort [18], access to these resources is often restricted to trusted research environments due to anonymisation protocols. Moreover, much of the publicly available data lacks critical demographic and ethnic metadata, with estimates suggesting that approximately 74% of these datasets do not include such information. As a result, even healthcare providers with extensive retinal image collections, ranging from ten thousand to ten million scans, face challenges in generating truly representative [19], integrative datasets.
@@
More broadly, a substantial gap exists in integrating retinal imaging data with molecular, clinical, and diverse datasets. Such integration is crucial for capturing the multi-dimensional nature of complex diseases [20]. However, epidemiological models, such as cohort, longitudinal and population-based studies, have successfully mapped disease trajectories in other domains [21]; their application to oculomics remains onerous [4]. The combined burden of global communicable and non-communicable diseases has further highlighted the need for such inclusive, multi-modal datasets that account for diversity across ethnicity, age, and geography [22].
@@
These shortcomings highlight the urgent need for a next-generation framework to unify fragmented data sources and enable deeper clinical insights. This includes robust epidemiological frameworks, large-scale imaging biobanks, and AI-enabled platforms harmonising disparate data types. Unified and multidimensional data frameworks—incorporating molecular signatures, imaging biomarkers, and longitudinal clinical profiles—hold the potential to provide a more holistic understanding of disease onset and progression. These frameworks, already successful in fields like oncology and cardiology [23], are both timely and necessary for their adoption in oculomics [5].
@@
As a result of these converging needs and opportunities, oculomics has emerged as a leading frontier in multimodal healthcare research, with biomarkers to infer both ocular and systemic health [24]. The retina is uniquely positioned for this purpose due to its rich vascular and neural architecture, which reflects broader physiological states throughout the body. As a transparent and non-invasive window into the vascular and neurobiological environment, RI allows early detection of systemic conditions. This makes the retina an ideal candidate for multimodal investigation in both research and clinical settings [25].
@@
Recent developments in automated image analysis tools have further enhanced retinal imaging capabilities. The AutoMorph software developed by Zhou et al. [1] enables a precise, large-scale extraction of morphological features from the retinal vasculature in a fully automated pipeline. This advancement significantly supports high-throughput oculomics research by offering standardised and scalable morphological analysis of retinal structures.
@@
In recent years, oculomics has emerged as a highly interdisciplinary domain, yet there remains a lack of cohesive literature synthesising its technological, clinical, and methodological advancements. Most existing reviews focus either on specific imaging techniques or isolated systemic diseases, without bridging the broader implications of AI-powered retinal analysis for integrated healthcare [26, 27, 6, 28, 29]. Given the rapid development of retinal image processing tools, dataset accessibility, and AI-driven diagnostic models, a comprehensive review is now both necessary and timely. This survey aims to fill that gap by contextualising classical foundations, highlighting cutting-edge innovations like AutoMorph, and proposing a unified framework for future research. This survey paper presents a structured synthesis of advancements in retinal imaging, AI integration, and multi-modal health data. It highlights key breakthroughs, identifies current gaps in clinical translation, and offers a roadmap for future interdisciplinary collaborations in oculomics. Through this, we aim to support the development of robust, explainable, and scalable tools capable of revolutionising healthcare delivery across disciplines.
@@ II-A Ocular Anatomy and Physiology
The field of ophthalmology has undergone significant evolution, marking a remarkable advancement in understanding, diagnosing, and treating eye diseases over the centuries. This evolution is driven by the imperative to comprehend a variety of ocular conditions arising from the diverse cellular, tissue, and neural components present within the eye. Each of these components is intricately linked to specific diseases, highlighting the necessity of studying the eye’s anatomy for improved diagnostic and early intervention strategies. Among these anatomical structures, externally visible components such as the cornea, iris, and pupil are essential for regulating light entry. In contrast, internal structures, namely the lens, retina, and optic nerve, play critical roles in visual processing and are central to clinical imaging and disease detection.
The visible components of the human eye include the cornea, iris, and pupil portrayed in Figure 1. The cornea, a transparent, curved structure, serves as the outermost layer of the eye, while the iris, the colored part of the eye, may manifest in various hues such as blue, green, or brown. Central to these features is the pupil, which appears black. Through fundoscopy, retinal images reveal deeper anatomical structures, as shown in Figure 2 (b). The cornea functions predominantly as a refractive lens, contributing approximately three-fourths of the eye’s focusing power, with the remaining one-fourth attributed to the actual lens. Light passes through the pupil and the lens before reaching the retina, traversing the vitreous humour, a gel-like substance that maintains ocular shape, supports the retina, and facilitates proper optical functioning.
The retina, situated at the posterior aspect of the eye, is a critical structure sensitive to light. Cross-sectional analysis of the retina uncovers six distinct classes of neurons (Figure 2 (b) ): cones, rods, amacrine cells, horizontal cells, bipolar cells, and ganglion cells. Each neuron type performs specific functions, yet collectively contributes to the overall process of vision. The retina is further encased by the choroid, illustrated in Figure 2 (a), with the optic nerve transmitting visual information to the brain’s visual cortex, culminating in the perception of sight. Notably, the macula functions as the eye’s focal point, wherein central vision is processed [30, 31].
Although the retina is often described as opaque, it is dependent on the support of the retinal pigment epithelium, choroid, and sclera. The choroid and retinal vasculature supply approximately 65% and 35% of the blood flow to the retina, respectively. A plethora of significant diseases can manifest within the retina, with origins either localised to the eye or connected to systemic conditions. As disorders affecting any of these anatomical components can lead to substantial visual impairment, a profound understanding of ocular anatomy and physiology is essential for effective diagnosis and treatment [32]. This section discusses several diseases, with additional conditions to be explored in forthcoming sections, addressing various aspects pertinent to ocular health.
Figure 1: External anatomy of the human eye, showing the cornea, iris, and pupil. These structures regulate how light enters the eye and initiate the visual processing pathway
Figure 2: Detailed cross-sectional illustration of the human retina and associated neuronal layers. Image reproduced from [32]

### v2

@@ Abstract
The unique vascularized anatomy of the human eye, encased in the retina, provides an opportunity to act as a window for human health. The retinal structure assists in assessing the early detection, monitoring of disease progression and intervention for both ocular and non-ocular diseases. The advancement in imaging technology leveraging Artificial Intelligence has seized this opportunity to bridge the gap between the eye and human health. This approach facilitates the unveiling of systemic health insights from the ocular system and the identification of non-invasive surrogate markers for timely intervention and detection. The new frontiers of oculomics in ophthalmology cover both ocular and systemic diseases, and are getting more attention to explore them. In this survey paper, we explore the evolution of retinal imaging techniques, the dire need for the integration of AI-driven analysis, and the shift of retinal imaging from classical techniques to oculomics. We also discuss some hurdles that may be faced in the progression of oculomics, highlighting the research gaps and future directions.
@@ I Introduction
The retina, the crucial light-sensitive tissue at the back of the eye, is a treasure trove of insights for ophthalmologists, clinicians, and technicians. It plays a pivotal role in the early prediction and intervention of both ocular and non-ocular disorders. Retinal imaging, a key interpretation, encompasses a variety of techniques designed to visualise and analyse the retina Zhou et al. (2022). Colour Fundus photography (CFP) captures detailed, subtle changes in the ocular system, while optical coherence tomography (OCT) provides a cross-sectional view, and OCT angiography visualises blood flow. These modalities, serving as vital diagnostic tools, enable healthcare professionals to examine the intricate structure of the retina, including the macula, optic disc, and retinal vasculature. This non-invasive assessment allows for the early detection and monitoring of ocular diseases like diabetic retinopathy, glaucoma, macular degeneration, and other related diseases that can lead to significant vision impairments if left unattended Ringel et al. (2021).
@@
Beyond ocular health, retinal imaging also provides insight into detecting systemic diseases. The retina’s unique accessibility and the visibility of its microvasculature make it a valuable indicator of overall health. For example, changes observed in the blood vessels and layers of the retina can trigger cardiovascular Zekavat et al. (2022), neurodegenerative Bsteh et al. (2024), and cerebrovascular diseases Cheung et al. (2013). That is why the eye serves as a window of overall health in the association of ocular and systemic health Cheung et al. (2012). High-resolution imaging can further reveal fine-grained features and act as a facilitator for timely intervention and personalized care. This contributes to improved patient outcomes and a better understanding of the interrelation between bodily systems Honavar (2022).
@@
Oculomics studies ocular biomarkers, using advanced imaging technologies to probe the eye in relation to other organ systems Wagner et al. (2020) through shared physiological functions, embryological origins, and anatomical structures Cheung et al. (2012). It aims to establish correlations between these markers and systemic diseases Honavar (2022). Growing evidence indicates that oculomics is significant in identifying clinical manifestations in ocular markers, such as retinal vessel patterns and neural layer thickness, which can reveal the presence of systemic diseases such as diabetes, cardiovascular diseases (CVDs), neurological issues, and stress-related disorders Hanssen et al. (2022). This field is driven by three significant advances, including the widespread availability of high-resolution ophthalmic imaging, the emergence of big data, and the development of automated artificial intelligence (AI) software. Together, it has the potential to revolutionise disease management by providing a comprehensive and accessible approach to monitoring and maintaining human health Zhu et al. (2025).
@@
AI is at its peak in the era of big data and parallel processing. The AI model is based on recorded data, and disciplines with data can easily participate in it. In the current epoch, it is hard to conclude that there may be fields in academia and industry that can thrive without data. Therefore, no discipline will be empty-handed when discussing data. AI models are now prevalent and deployed in every aspect of life Lyu et al. (2021). Both intrinsic and extrinsic deployments have positively impacted our society, especially in socioeconomic and socio-health aspects. Renowned computer scientist Dr. Knuth revealed that biology has at least 500 years of exciting problems to explore Bansal et al. (2023). This perspective encourages the computing and biomedical research communities to address challenges in life science collaboratively.
@@
AI is transforming retinal imaging and ocular studies by automating and enhancing the analysis of complex visual data. AI algorithms, especially deep learning (DL), excel in identifying subtle patterns and abnormalities within retinal images that human observers might overlook Hassan et al. (2024). This automation significantly speeds up the screening process for ocular and non-ocular diseases, allowing for earlier detection and intervention. In addition, AI can enable the development of predictive models by evaluating the vast data set, encompassing the assessment of the individual patient’s risk factors and anticipating disease progression, paving the way for personalised medicine in ophthalmology Ting et al. (2019); Schmidt-Erfurth et al. (2018).
@@
Eye diagnostic research is shifting from its classical origin to the current AI software tools. Now, it’s beyond the classical methods, such as preprocessing, model training, and limited clinically relevant output. These limited circumstances hinder the quantification of retinal structures at the minute and micro levels to address the objective measurements, which are crucial for the efficacy of effective treatments and the monitoring of disease progression in the growing field of oculomics. Likewise, for meaningful research, high-quality and large health datasets comprising rich annotated data are crucial, which advances the role of statistical, computational, and AI-driven techniques Brown et al. (2024); Zhou et al. (2023). Furthermore, this framework can be further nourished by integrating retinal imaging with clinical information, namely electronic health records and genetic data, including diverse demographic and ethnic data. However, translating this vision into practice remains a considerable challenge. Existing datasets often lack the necessary integration across epidemiological, clinical, and imaging domains, limiting the development of a unified infrastructure supporting oculomics research and clinical application Wagner et al. (2020); De Fauw et al. (2018); Poplin et al. (2018).
@@
Despite the increasing availability of large-scale datasets, significant limitations impede their practical use in oculomics. Even though there are several Retinal imaging repositories, including those of Moorfields Eye Hospital De Fauw et al. (2017), UK Biobank Collins (2012), and the 10k Cohort Shilo et al. (2021), access to these resources is often restricted to trusted research environments due to anonymisation protocols. Moreover, much of the publicly available data lacks critical demographic and ethnic metadata, with estimates suggesting that approximately 74% of these datasets do not include such information. As a result, even healthcare providers with extensive retinal image collections, ranging from ten thousand to ten million scans, face challenges in generating truly representative Khan, Liu, Nath, Korot, Faes, Wagner, Keane, Sebire, Burton & Denniston (2021), integrative datasets.
@@
More broadly, a substantial gap exists in integrating retinal imaging data with molecular, clinical, and diverse datasets. Such integration is crucial for capturing the multi-dimensional nature of complex diseases Agarwal et al. (2024). However, epidemiological models, such as cohort, longitudinal and population-based studies, have successfully mapped disease trajectories in other domains Pérez-Guerrero et al. (2024); their application to oculomics remains onerous Wagner et al. (2020). The combined burden of global communicable and non-communicable diseases has further highlighted the need for such inclusive, multi-modal datasets that account for diversity across ethnicity, age, and geography Vos et al. (2020).
@@
These shortcomings highlight the urgent need for a next-generation framework to unify fragmented data sources and enable deeper clinical insights. This includes robust epidemiological frameworks, large-scale imaging biobanks, and AI-enabled platforms harmonising disparate data types. Unified and multidimensional data frameworks, including molecular signatures, imaging biomarkers, and longitudinal clinical profiles, have the potential to provide a more holistic understanding of disease onset and progression. These frameworks, already successful in fields such as oncology and cardiology Wang, Liu, Li, Zhang, Sun, Wang, Li, Li, Li, He et al. (2024), are timely and necessary for their adoption in oculomics Hanssen et al. (2022).
@@
As a result of these converging needs and opportunities, oculomics has emerged as a leading frontier in multimodal healthcare research, with biomarkers to infer both ocular and systemic health Pearson et al. (2022). The retina is uniquely positioned for this purpose due to its rich vascular and neural architecture, which reflects broader physiological states throughout the body. As a transparent and non-invasive window into the vascular and neurobiological environment, retinal imaging allows early detection of systemic conditions. This makes the retina an ideal candidate for multimodal investigation in both research and clinical settings Ran et al. (2024).
@@
Recent developments in automated image analysis tools have further enhanced retinal imaging capabilities. The AutoMorph software developed by Zhou et al. Zhou et al. (2022) enables a precise, large-scale extraction of morphological features from the retinal vasculature in a fully automated pipeline. This advancement significantly supports high-throughput oculomics research by offering standardised and scalable morphological analysis of retinal structures.
@@
In recent years, oculomics has emerged as a highly interdisciplinary domain, yet there remains a lack of cohesive literature synthesising its technological, clinical, and methodological advancements. Most existing reviews focus either on specific imaging techniques or isolated systemic diseases, without bridging the broader implications of AI-powered retinal analysis for integrated healthcare Jin & Ye (2022); Akpinar et al. (2024); Zhu et al. (2025); Lee et al. (2024); Li, Yin, Wang, Wang, Qiang & Jiang (2024). Given the rapid development of retinal image processing tools, dataset accessibility, and AI-driven diagnostic models, a comprehensive review is now both necessary and timely. This survey aims to fill that gap by contextualising classical foundations, highlighting cutting-edge innovations like AutoMorph, and proposing a unified framework for future research. This survey paper presents a structured synthesis of advancements in retinal imaging, AI integration, and multi-modal health data. It highlights key breakthroughs, identifies current gaps in clinical translation, and offers a roadmap for future interdisciplinary collaborations in oculomics. Through this, we aim to support the development of robust, explainable, and scalable tools capable of revolutionising healthcare delivery across disciplines.
@@ II-A Ocular Anatomy and Physiology
The eye is a living optical system, and its role as a window into systemic health is grounded in its unique anatomy and physiology, where refractive, neuronal, and vascular components interact to enable vision and reflect disease processes. Each element of this system, illustrated in Figure 1, is not only vital for sight, but is also sensitive to disturbance, which means that localised abnormalities often signal both ocular pathology and broader systemic disease.
At the outer surface, the sclera, iris, and pupil regulate the entry of light and protect the delicate internal structures, as portrayed in Figure 1a. The cornea, while anatomically part of the external coat of the eye, is transparent and therefore not readily visible in a superficial view. In cross-sectional orientation (Figure 1b), the cornea is seen as a curved dome that provides nearly three-quarters of the eye’s total refractive power. Its highly ordered collagen fibrils maintain curvature and transparency Abràmoff et al. (2010), but when disrupted, as in keratoconus or corneal scarring, irregular astigmatism and impaired vision result Kreps et al. (2021). Surrounding it, the sclera forms the opaque, fibrous shell that preserves the globe’s shape and stability. Excessive thinning and stretching of the sclera underlie high myopia, now a growing global public health concern Jonas et al. (2019). The iris adjusts the pupil aperture to balance light entry: it widens in low-light environments and constricts in bright conditions, safeguarding the retina from excess light exposure.
Behind the iris lies the crystalline lens, a biconvex structure suspended by zonular fibers. The lens fine-tunes focus through accommodation, altering its curvature to shift between near and distant vision. With aging, the lens gradually hardens and its proteins aggregate, producing cataract, the world’s leading cause of reversible blindness Wishart et al. (2021). Light then traverses the vitreous humor, a transparent gel that fills the posterior chamber. This medium stabilizes the ocular globe and transmits light without distortion, but it is not static. Over time, the vitreous liquefies and contracts, often detaching from the retina. Such posterior vitreous detachment can predispose to retinal tears and even retinal detachment, both of which pose a serious risk to sight.
The retina lies at the back of the eye and functions as the sensory core of vision. It is a multilayered neural tissue where photons are converted into electrical signals. As shown in Figure 1c, the retinal microstructure consists of photoreceptors (rods and cones), bipolar, horizontal, and amacrine cells, ganglion cells, and the retinal pigment epithelium.Specialised photoreceptors, rods for dim-light and peripheral vision, and cones for colour perception and fine detail, initiate this process. Their output is integrated by horizontal, bipolar, and amacrine cells, which refine contrast and spatial information. Finally, ganglion cells collect these signals, and their axons converge to form the optic nerve, transmitting information to the brain’s visual cortex. Within the retina, the macula supports high-acuity central vision, while the peripheral retina detects movement and supports night vision Cholkar et al. (2013); Liang et al. (2023). Cross-sectional analyses, such as those provided by OCT, reveal this layered architecture in vivo and demonstrate how dysfunction at any level, whether photoreceptor loss in age-related macular degeneration (AMD) or microvascular leakage in DR, directly alters visual capacity. Figure 1b highlights these internal structures in cross-section, including the cornea, lens, vitreous humor, retina, choroid, and optic nerve
The retina is sustained by an intricate network of supporting tissues and circulations. The retinal pigment epithelium (RPE) nourishes photoreceptors, regulates waste clearance, and recycles visual pigments. The choroid, a dense vascular layer, supplies approximately 65% of retinal oxygen and nutrients, while the retinal vasculature provides the remaining 35% Abràmoff et al. (2010). This dual circulation reflects the enormous metabolic demand for the retina. Because retinal vessels are optically accessible, their geometry, summarised by traits such as vessel caliber, fractal dimension, tortuosity, and vessel density, provides quantifiable endophenotypes. These vascular signatures have been linked not only to local conditions such as DR but also to systemic disease, including hypertension, cardiovascular disease, stroke, and even renal dysfunction Cheung et al. (2013); Ikram et al. (2013)
The optic nerve forms the final bridge between the eye and the brain. Its vulnerability is twofold: localised ocular pathologies, such as glaucoma, that progressively damage retinal ganglion cell axons, while systemic neurodegeneration also manifests here. The thinning of the retinal nerve fiber layer (RNFL) has been observed in both Alzheimer’s and Parkinson’s disease Snyder et al. (2021), suggesting that the eye can act as a biomarker of central nervous system health. This role, both in transmitting visual signals and in reflecting neurological status, cements its place in the investigation of systemic disease.
Collectively, the eye is not only an organ of sight, but also a mirror of health. Each component, from the cornea that bends light to the retina that transduces it into neural signals to the optic nerve that delivers these signals to the brain, offers unique diagnostic insight. Because these structures are accessible through modern imaging modalities, their disruption reveals both ocular disease and systemic pathology. This foundational understanding of anatomy and physiology has historically guided eye research and now underpins the transition to advanced imaging and oculomics, which will be explored in the following section on the historical development of ophthalmology.
Figure 1: Structural organization of the human eye at multiple anatomical levels. (a) External anatomy of the eye showing the pupil (central opening), iris (colored ring), and sclera (white outer coat that maintains ocular shape and protection). (b) Cross-sectional schematic illustrating the major internal structures including the cornea, iris, lens, vitreous humor, sclera, choroid, retina, blood vessels, and optic nerve. This orientation highlights the key optical and supportive components of the eye. (c) Retinal microstructure depicting the layered neuronal architecture, including photoreceptors (rods and cones), bipolar, horizontal, and amacrine cells, ganglion cells, and the pigment epithelium.